# NeuroPilot: A Realtime Brain-Computer Interface system to enhance concentration of students in online learning

### v1

@@ III-A2 Filtering
The EEG signals have been filtered using a bidirectional Butterworth bandpass with a cutoff frequency of 0.5–64 Hz as a standard range in this research field with an optimal order of 3 found through graphical comparisons. The problem arises due to edge effects caused by filter initialization and group delay as seen in Fig. 3a. To overcome this, 250 samples has been trimmed from both ends of each segment to get clean signal with almost no phase distortion after the filtering is done. The procedure can be visualized from Fig. 3b. This method adds a constraint to the selection of segment length for experimentation which is $L>500$ samples. This means that the new segments found after trimming will have a length of (L500L-500) samples which gives another constraint of setting the ratio $r$. The constraint can be understood as $r\times L>500$. The logical explanation can be understood from the overall segmentation and trimming procedure as given in Fig. 3c where the sequential overlapping and trimming of the segments are shown along with the obtained segments. Then the signal is passed through a notch filter of 50 Hz to avoid power line interference.
@@ III-A3 Artifact Removal
Non-invasive EEG signals are often contaminated by artifacts, including eye blinks. Two steps have been followed in this research for noise removal. First, the filtered segments are checked for high-amplitude values greater than a threshold of 150 $\mu\text{V}$ as meaningful EEG data lie below this value [38]. If any segment contains at least one high amplitude value, then that segment is dropped from the dataset. Correlation and interpolation are commonly used for high-amplitude portions [39, 40], but the segment rejection approach has been used here to exclude artificial continuations, especially given the abundance of collected data. Following this, eye blinks are removed through Ensemble Empirical Mode Decomposition (EEMD) [41]. Peak indices are extracted from the filtered EEG signal using the find_peaks function from the SciPy library [42]. Then three conditions have been imposed to taking the region around the peak. If a peak index falls within the starting 40 or ending 80 values of the segment, then the starting 120 or ending 120 values are taken for EEMD. However, if the peak index falls within the rest of the region then the 40 values before and 80 values after the peak index are taken. These indices are taken from rigorous experimentation to get optimal artifact removal. The signal found after doing EEMD is passed through an uniform filter to smooth residual high-frequency noise.
@@ III-B Feature Extraction
Following pre-processing, feature extraction is applied to the segments. A comprehensive set of 458 features have been extracted from the EEG signals to capture time-domain, frequency-domain, wavelet-domain, and statistical characteristics. Time-domain features included commonly used descriptors such as Hjorth parameters (e.g., complexity, activity), median (md), normalized second difference (n2d) etc. Frequency-domain features encompassed spectral moments (e.g., centroid, spread), entropy-based measures, and spectral shape descriptors (e.g., crest, flatness, roll-off). Additionally, wavelet packet transform (WPT) has been used to extract sub-band energy and entropy features which captures localized time–frequency dynamics. These features have been selected based on their proven effectiveness in attention evaluation through EEG and biosignal analysis. Eight additional features obtained from the Crimson SDK have been appended to the initial 458 features, resulting in a total of 466 features. These include six frequency band powers of the raw EEG signal, along with attention and meditation scores for each segment. As these features are readily available through the open-source Crimson SDK and have been commercially utilized, they are included in the feature set to explore their potential contribution to the classification task. After feature extraction, all computed feature matrices are examined for invalid values. Such invalid values can arise from divisions by zero or logarithmic transformations of zero-valued data, particularly in spectral features. Any features containing NaN (Not a Number) or infinite (Inf) values are identified and replaced with zero. Zero has been used as it preserves dimensionality, enables batch processing, and acts as a neutral, non-informative value without introducing imputation bias. This step ensures numerical stability and prevents errors in subsequent processing and model training.
@@ III-C Feature Selection
The next step after feature extraction is selection of most relevant features for training the data. This is necessary to ensure that the final model is trained on the best set of features that contributes to the classification performance. The selection has been done in two steps: Pearson correlation filtering (PCF), followed by Recursive Feature Elimination (RFE) using Support Vector Machine (SVM).
@@ III-C1 Pearson Correlation Filtering
PCF has been used to reduce high dimensionality and optimize computational efficiency which works through selecting one from each highly correlated pair when the coefficient of correlation exceeds a threshold $p$. The value of $p=0.8$ (details in section III.D) has been used to remove the highly correlated features, leaving 108 in the set for further analysis. The Pearson correlation coefficient is defined as:
@@ III-C2 Recursive Feature Elimination with SVM
Subsequently, RFE with SVM has been exploited to find the meaningful set of features required for optimal model training. It is a supervised feature selection method that removes the least important features recursively based on the weights assigned by a linear SVM model. At each iteration, the model is trained on the current feature set, and features with the smallest absolute weights (least contribution to the decision function) are eliminated. This process continues until the desired number of features $d=50$ (details in section III.D) is retained.
First, RFE has been applied for each subject (identified by user_id) using the linear SVM. The regularization parameter $C_{1}$ for this SVM has been iterated through a range (0.01 to 100), to account for varied model complexity and feature importance. For each $C_{1}$ value, a set of 50 top-ranked features is obtained. These sets are not directly finalized but evaluated through another SVM with radial basis function (RBF) kernel. The classification performance of this secondary model serves as the basis for selecting the optimal feature set. This subject-focused process has been designed so that the selected features are not only ranked high based on linear separability but also demonstrate better generalization of the model. Finally, a set of 9 common features is obtained through finding the most frequently selected features across all subjects. Out of these features, the relative powers of alpha and delta band, ratio of low beta to summation of alpha and theta band, average power of low beta band, and median of high beta band show strong correlations with those reported in the literature as contributing to sustained attention [19, 18, 20, 21]. These consensus-based features has been used for final model training. The selected features for final training along with their meanings are given in Table II.
@@ III-D Hyperparameter Tuning
All parameters defined from pre-processing to feature selection have been systematically varied to evaluate their impact on classification performance through rigorous assessments. For segment length L, values ranging from 5 to 8 paired with overlap ratios ranging from 0.3 to 0.7 have been used for analysis. The pair of $(L,r)=(1750,0.7)$ has been found to be optimal through comparisons of model performance across all tested values. Similar approach has been followed for Pearson correlation coefficient threshold $P$ (0.6 to 0.95 in increments of 0.05) and optimal number of features $d$ (30 to 80 in increments of 10). The best trade-off between model performance and generalization has been found for $P=0.8$ $d=50$. For RFE, the linear SVM has been trained on $C_{1}\in\{0.01,\ 0.1,\ 1,\ 10,\ 100\}$ and the secondary SVM for identification of optimal feature set has been trained on a nested grid search of two parameters $C_{2}\in\{0.01,\ 0.1,\ 1,\ 10,\ 100\}$ and $\gamma\in\{0.001,\ 0.01,\ 0.1,\ 0.5,\ 1,\ 10\}$. The combination yielding the highest performance determines the best feature set corresponding to the optimal $C_{1}$. Final classification is also tuned using a separate grid search on the same range of $C_{2}$ and $\gamma$ which is applied on the common features obtained across all users. The optimal pair of $(C_{2},\gamma)=(0.01,0.5)$ has been found to yield the best classification performance.
@@ III-E Classification
SVMs have been widely adopted for EEG classification tasks due to their robustness in high-dimensional spaces, effectiveness with small sample sizes, and ability to articulate non-linear decision boundaries through kernel functions. Thus, an SVM with RBF kernel has been selected for capturing non-linear relationships in the feature space. The SVM finds a hyperplane that maximizes the margin between classes while allowing some misclassifications which is controlled by the regularization parameter $C_{2}$}. The kernel coefficient $\gamma$ determines the influence of individual training samples on the decision boundary, shaping how tightly the model fits local patterns in the feature space.
The training pipeline as explained in Algorithm 1 includes fitting the model on a nested grid search of $C_{2}$ and $\gamma$. The evaluation is performed through leave-one-subject-out (LOSO) cross-validation. LOSO is an appropriate choice to effectively assess the model’s generalizability on unseen subjects which closely reflects real-world application scenarios.
Input: Dataset $\mathcal{D}$ with subject labels user_id, hyperparameter grids $\mathcal{C}_{2}$, $\Gamma$
Output: Optimal hyperparameters per subject and performance metrics
1
2Obtain set of unique subjects $U$ from $\mathcal{D}$ ;
3 foreach subject $u_{\text{test}}\in U$ do
4 $\mathcal{D}_{\text{test}}\leftarrow$ data from $u_{\text{test}}$ ;
5 $\mathcal{D}_{\text{train}}\leftarrow\mathcal{D}\setminus\mathcal{D}_{\text{test}}$ ;
6
 // Nested grid search on training data
7   Initialize $\texttt{best\_score}\leftarrow 0$ ;
8 foreach $C_{2}\in\mathcal{C}_{2}$ do
9    foreach $\gamma\in\Gamma$ do
10         Train SVM with RBF kernel using $(C_{2},\gamma)$ on $\mathcal{D}_{\text{train}}$ ;
11         Evaluate on $\mathcal{D}_{\text{test}}$ to get score ;
12       if score >> best_score then
13          $\texttt{best\_score}\leftarrow$ score ;
14            Save $(C_{2},\gamma)$ as optimal for $u_{\text{test}}$ ;
15
16
17
18  Store best score and best $(C_{2},\gamma)$ for $u_{\text{test}}$ ;
19
20Compute average performance across all subjects ;
Algorithm 1 Model Training And Validation Using LOSO
After a rigorous grid search, the pairs of $C_{2}$ and $\gamma$ yielding the best classification performance for the subjects are recorded. From these pairs, the most frequent one is selected for fitting the final model which is then used for real-time evaluation.
@@ III-F Evaluation Metrics
The performance of the classifier has been evaluated using standard metrics including accuracy, precision, recall, and F1-score. As the dataset has been balanced per class during segmentation, these four metrics are sufficient for evaluating the model performance.
@@ IV Experimental Results
This section provides the performance evaluation of the trained model and the ablation studies that have led to its optimized version.
@@ IV-A Cross-Validation Results
The model performance has been analyzed using the LOSO cross-validation procedure (discussed in section III.E). The optimized model performance on each fold (fold implying the data of each subject as test data while taking all the others as train data) has been shown in Fig. 4. The optimized model achieves the average test accuracy of 88.77%. Additionally, it attains a precision of 90.49%, recall of 88.77%, and F1 score of 89.03% which indicates strong predictive capability with a good balance between correctly identifying positive cases and minimizing false positives.
@@ IV-B Model Ablation Studies
Initially, 458 features excluding the 8 features provided by the Crimson SDK, are used to train the SVM which gave an accuracy of 76.44%. Subsequently, the 8 SDK features are incorporated, and the model is retrained, achieving the current best performance. Among these additional features, the attention and meditation scores are included in the final feature set. To assess whether the model’s performance is overly reliant on these two scores, it has been compared against the baseline performance obtained using only the Crimson SDK’s scores as the feature set. Crimson SDK has been selected as the baseline as it gives the scores which are associated with cognitive load. To have the accuracy from those scores as predictions to the EEG signals,two methods have been employed. The first method includes training the model on the two features and evaluating the metrics, which gave an accuracy of 84.13%. The second method involves the ratio of attention to medition score $r_{am}$ as the condition to classify the data. For all the n segments found, the predictions have been taken to be attentive if $r_{am}>1$ and non-attentive otherwise. The accuracy is found to be 82.27%. This shows that the proposed methodology outperforms the current interface of the FocusCalm headband modified for attention classification in online learning environments (details discussed in section V). A brief summary has been given in Table III for a clear overview.
@@ V Discussion
The overall research methodology differs from baseline methods (given the limited prior research addressing this objective within this specific environment). Therefore, the discussion focuses on the unique aspects of the approach, the rationale for their selection, their implementation, and the specific methods they outperform. Also, the limitations of the current research along with logical representations have been mentioned.
@@
The proposed EEG-based attention classification achieves 6.5% better accuracy than the Crimson SDK, which suggests that the combination of wavelet packet decomposition along with entropy features are effective in capturing attention state dynamics. Although baseline methods achieve satisfactory results, the data collection strategies of other baseline studies such as Rehman et al. and Huang et al. are based on instructions and there remained a room for data validation strategies which have been implemented in this research. The steps followed in this research ensure correct predictions which pertains to ground truths and this is crucial for real-time implementation.
@@
Initially, device selection was emphasized, as most previous research has relied on non-consumer-grade devices. For this study, the FocusCalm headband has been chosen for its portability, wearability, and ease of hands-on use, despite the potential for experiments with the Muse headband, which offers greater spatial coverage in real-time applications. Given our objective of developing a state-of-the-art real-time binary classification system, the FocusCalm device has proven sufficient for our purposes. The data collection protocol was designed to reflect the environmental setting of classrooms, with video stimuli consisting of lecture content prepared for students. The data validation process was established under the principle that only ground-truth labels are required. Instances can occur where a participant experiences cognitive load yet fails to answer a question. However, this does not affect labeling, as the protocol ensures that all selected data correspond to verified ground truths.
@@
The methods used in this research were selected through extensive trials of different approach combinations. For instance, both causal forward filtering and zero-phase second-order section (SOS) filtering were tested. While the latter removed phase distortion more effectively, it performed poorly at the start and end of segments during segment-wise processing, as it requires future data to operate. To address this, those edge portions were removed, ensuring both the absence of distortion and the independence from future data. Furthermore, to prevent data loss, the segmentation process used an overlapping ratio that retained common portions between consecutive segments. For classification, SVM was initially selected due to its proven effectiveness in EEG signal analysis. Conceptually, SVM can be viewed as a mechanism that iteratively identifies the most discriminative features in a high-dimensional space, analogous to how traditional EEG analyses iteratively isolate regions or frequency bands of interest to distinguish cognitive states. An initial assessment using various machine learning models was conducted, and the SVM achieved the highest performance, which is consistent with observations reported in the literature.
@@
After validating the data using LOSO, the optimal $(C_{2},\gamma)$ values were determined in two ways. The first method identified the most common $(C_{2},\gamma)$ pair that achieved the best metrics across all subjects. The second method selected the $(C_{2},\gamma)$ pair that yielded the highest metrics, regardless of its frequency among subjects. The mode pair, $(C_{2},\gamma)=(0.01,0.5)$, produced an average accuracy of 88.77%, whereas $(C_{2},\gamma)=(0.01,0.1)$ achieved a slightly higher average accuracy of 89.63%. Despite the higher accuracy, the former pair was adopted for real-time testing due to its more frequent favorable performance, which makes the model generalized and objective. Another key finding from the LOSO validation results was the model’s relatively poor performance on two subjects when taken as test set, although it has achieved satisfactory classification results for the other subjects as shown in Fig. 4. This indicates that the model can be made more robust by incorporating additional data.
VI Realtime Setup
VI-A Data-handling Procedure
The real-time implementation has been done through building a software interface in python. The raw EEG data stream from the headband device is handled using a circular buffer mechanism to ensure continuous real-time processing without data loss or overlap. A fixed-size buffer of capacity $B_{C}=10000$ is allocated to store incoming EEG samples, along with parallel buffers to record the corresponding attention and meditation metrics. This capacity is chosen to accommodate multiple segments while maintaining an effective trade-off between memory utilization and processing efficiency. Incoming EEG data packets are appended to the buffer. When the insertion point approaches the end of the buffer, the packet is split and wrapped around to the buffer’s beginning, ensuring efficient circular memory usage. This approach prevents overwriting unread data and maintains temporal continuity. A read pointer tracks the start position of data segments ready for analysis. Data availability is monitored by a counter that tracks the total unread samples in the buffer. Once the accumulated unread data reaches or exceeds a predefined window size of 1750 samples (length of segment), the application extracts fixed-length EEG segments for processing. After each window prediction, the read pointer advances a step size of 525 samples (to maintain the same overlap ratio of 0.7 for producing segments of training data), meaning that the window size is filled and processed every 2.1 seconds approximately. Segments are extracted with potential wrap-around handling which implies that, if the segment spans the end of the buffer, it is reconstructed by concatenating the tail portion with the head of the buffer to maintain continuity. In the processing function, the noise removal, feature extraction and classification using the trained model is done. Following this, the read pointer advances by a defined step size, and the unread data count decreases accordingly, supporting overlapping window analysis and ensuring no data gaps.
@@ VI-A Data-handling Procedure
Figure 5: Subfigures (a) and (b) represents the graphical user interface and the real-time working pipeline. Subfigure (c) denotes the changes between attentive and non-attentive states during the real-time system test on five participants.

### v2

@@ III-A2 Filtering
The EEG signals were filtered using a bidirectional Butterworth bandpass with a cutoff frequency of 0.5–64 Hz, which is a standard range in this research field, and an optimal order of 3 found through graphical comparisons. The problem arises due to edge effects caused by filter initialization and group delay as seen in Fig. 3a. To overcome this, 250 samples were trimmed from both ends of each segment to get clean signal with minimal phase distortion after filtering. The procedure is illustrated in Fig. 3b. This method added a constraint to the selection of segment length for experimentation which is $L>500$ samples. This means that the new segments found after trimming had a length of (L500L-500) samples, which gave rise to another constraint of setting the ratio $r$. The constraint can be understood as $r\times L>500$. The logical explanation is understandable from the overall segmentation and trimming procedure, depicted in Fig. 3c, where the sequential overlapping and trimming of the segments are shown along with the obtained segments. Finally, the signal was passed through a 50 Hz notch filter to remove power line interference.
@@ III-A3 Artifact Removal
Non-invasive EEG signals are often contaminated by artifacts, including eye blinks. Two steps were performed for noise removal. First, the filtered segments were checked for high-amplitude values greater than a threshold of 150 $\mu\text{V}$, as meaningful EEG data lie below this value [45]. If any segment contained at least one high amplitude value, then that segment was dropped from the dataset. Correlation and interpolation are commonly used for high-amplitude portions [46, 47], but the segment rejection approach was used to exclude artificial continuations, given the abundance of collected data. Following this, eye blinks were removed through Ensemble Empirical Mode Decomposition (EEMD) [48]. Peak indices were extracted from the filtered EEG signal using the find_peaks function from the SciPy library [49], which finds all the local maxima. Then three conditions were imposed to take the region around the peak. If a peak index fell within the starting 40 or ending 80 values of the segment, then the starting 120 or ending 120 values were taken for EEMD. However, if the peak index fell within the rest of the region then the 40 values before and 80 values after the peak index were taken. These indices were acquired through rigorous experimentation to get optimal artifact removal. The signal found after performing EEMD was then smoothed using a uniform filter to eliminate residual high-frequency noise.
@@ III-B Feature Extraction
Following pre-processing, feature extraction was applied to the segments. A comprehensive set of 458 features was extracted from the EEG signals to capture time-domain, frequency-domain, wavelet-domain, and statistical characteristics. Time-domain features included commonly used descriptors such as Hjorth parameters (e.g., complexity, activity), median (md), normalized second difference (n2d) etc. Frequency-domain features encompassed spectral moments (e.g., centroid, spread), entropy-based measures, and spectral shape descriptors (e.g., crest, flatness, roll-off). Additionally, wavelet packet transform (WPT) was used to extract sub-band energy and entropy features, which effectively captures localized time–frequency dynamics. These features were selected based on their proven effectiveness in attention evaluation through EEG and biosignal analysis. Eight additional features from the Crimson SDK (six frequency band powers along with attention and meditation scores) were appended to the initial 458 features, for a total of 466. These standard, open-source features were included to explore their potential contribution. Following feature extraction, a data cleaning step was performed. Any features containing NaN or Inf values (arise from divisions by zero or logarithmic transformations of zero-valued data) were replaced with zero. Zero was chosen as a neutral, non-informative value that preserves data dimensionality and ensures numerical stability for model training.
@@ III-C Feature Selection
The next step after feature extraction was selection of the most relevant features for training the data. This was necessary to ensure that the final model was trained on the best set of features that contribute to the classification performance. The selection was done in two steps: Pearson correlation filtering (PCF), followed by Recursive Feature Elimination (RFE) using Support Vector Machine (SVM).
@@ III-C1 Pearson Correlation Filtering
PCF was used to reduce high dimensionality and optimize computational efficiency, which works by selecting one from each highly correlated pair when the coefficient of correlation exceeds a threshold $p$. The value of $p=0.8$ (details in section III.D) was used to remove the highly correlated features, leaving 108 in the set for further analysis. The Pearson correlation coefficient is defined as:
@@ III-C2 Recursive Feature Elimination with SVM
Subsequently, RFE with SVM was applied to find the meaningful set of features required for optimal model training. It is a supervised feature selection method that removes the least important features recursively based on the weights assigned by a linear SVM model. At each iteration, the model was trained on the current feature set, and features with the smallest absolute weights (least contribution to the decision function) were eliminated. This process was continued until the desired number of features, $d=50$ (details in section III.D), was retained.
First, RFE was applied for each subject (identified by user_id) using a linear SVM. The regularization parameter $C_{1}$ for this SVMwas iterated over a range (0.01 to 100), to account for varied model complexity and feature importance. For each $C_{1}$ value, a set of 50 top-ranked features was obtained. These sets were then evaluated through another SVM with a radial basis function (RBF) kernel. Classification performance of this secondary model served as the basis for selecting the optimal feature set. This subject-focused process was designed so that the selected features were not only ranked high based on linear separability but also demonstrated better generalization of the model. Finally, a set of 9 common features was obtained by finding the most frequently selected features across all subjects. These included the relative powers of alpha and delta band, ratio of low beta to summation of alpha and theta band, average power of low beta band, and median of high beta band. These features showed strong correlations with those reported in the literature as contributing to sustained attention [23, 22, 24, 25]. These consensus-based features were used for final model training. The selected features for final training along with their meanings are given in Table II.
@@ III-D Hyperparameter Tuning
All parameters defined from pre-processing to feature selection were systematically varied to evaluate their impact on classification performance through rigorous assessments. For the segment length L, values ranging from 5 to 8, paired with overlap ratios ranging from 0.3 to 0.7, were used for analysis. The pair of $(L,r)=(1750,0.7)$ was found to be optimal through comparisons of model performance across all tested values. A similar approach was followed for the Pearson correlation coefficient threshold $P$ (0.6 to 0.95 in increments of 0.05) and the optimal number of features $d$ (30 to 80 in increments of 10). The best trade-off between model performance and generalization was found for $P=0.8$ and $d=50$. For RFE, the linear SVM was trained on $C_{1}\in\{0.01,\ 0.1,\ 1,\ 10,\ 100\}$, and the secondary SVM for identifying the optimal feature set was trained using a nested grid search over two parameters: $C_{2}\in\{0.01,\ 0.1,\ 1,\ 10,\ 100\}$ and $\gamma\in\{0.001,\ 0.01,\ 0.1,\ 0.5,\ 1,\ 10\}$. The combination yielding the highest performance determined the best feature set corresponding to the optimal $C_{1}$. Final classification was further tuned using a separate grid search on the same range of $C_{2}$ and $\gamma$, applied to the common features obtained across all users. The optimal pair $(C_{2},\gamma)=(0.01,0.5)$ was found to yield the best classification performance.
@@ III-E Classification
SVMs are widely adopted for EEG classification tasks due to their robustness in high-dimensional spaces, effectiveness with small sample sizes, and ability to model non-linear decision boundaries through kernel functions. Therefore, an SVM with an RBF kernel was selected to capture non-linear relationships in the feature space. The SVM finds a hyperplane that maximizes the margin between classes while allowing some misclassifications, which are controlled by the regularization parameter $C_{2}$}. The kernel coefficient $\gamma$ determines the influence of individual training samples on the decision boundary, thereby shaping how tightly the model fits local patterns in the feature space.
The training pipeline included fitting the model on a nested grid search of $C_{2}$ and $\gamma$. The evaluation was performed through leave-one-subject-out (LOSO) cross-validation. LOSO is an appropriate choice to effectively assess the model’s generalizability to unseen subjects, thereby closely reflecting real-world application scenarios.
After a rigorous grid search, the pairs of $C_{2}$ and $\gamma$ yielding the best classification performance for the subjects were recorded. From these pairs, the most frequent combination was selected to fit the final model, which was then used for real-time evaluation.
@@ III-F Evaluation Metrics
The performance of the classifier was evaluated using standard metrics, including accuracy, precision, recall, and F1-score. Since the dataset was balanced per class during segmentation, these metrics were sufficient to estimate the model performance.
@@ IV Experimental Results
This section presents the performance evaluation of the trained model and the ablation studies that led to its optimized version.
@@ IV-A Cross-Validation Results
The model performance was analyzed using the LOSO cross-validation procedure (discussed in section III.E). The optimized model’s performance on each fold (fold implying the data of one subject as test data while taking all the others as train data) is shown in Fig. 4. The optimized model achieved an average test accuracy of 88.77%. Additionally, it attained a precision of 90.49%, recall of 88.77%, and F1-score of 89.03%, which indicates strong predictive capability with a good balance between correctly identifying positive cases and minimizing false positives.
@@ IV-B Model Ablation Studies
Initially, 458 features excluding the 8 features provided by the Crimson SDK, were used to train the SVM, which gave an accuracy of 76.44%. Subsequently, the eight SDK features were incorporated, and the model was retrained, achieving the current best performance. Among these additional features, the attention and meditation scores were included in the final feature set. To assess whether the model’s performance is overly reliant on these two scores, it was compared against the baseline performance obtained using only the Crimson SDK’s scores as the feature set. The Crimson SDK was selected as the baseline, as it provides the scores associated with cognitive load. To evaluate the predictive accuracy of these scores for the EEG signals, two methods were employed. The first method included training the model on the two features and evaluating the metrics, which gave an accuracy of 84.13%. The second method involved the ratio of attention to meditation score $r_{am}$ as the condition to classify the data. For all the n segments found, the predictions were taken to be attentive if $r_{am}>1$ and non-attentive otherwise. The accuracy was found to be 82.27%. This shows that the proposed methodology outperforms the current interface of the FocusCalm headband modified for attention classification in online learning environments (details discussed in section V). A brief summary is provided in Table III for a clear overview.
@@ V Discussion
The overall research methodology differs from baseline methods (given the limited prior research addressing this objective within this specific environment). Therefore, the discussion focuses on the unique aspects of the approach, the rationale for their selection, their implementation, and the specific methods they outperform. Finally, this section addresses the limitations of the current research and provides a logical justification for the methods selected.
@@
The proposed EEG-based attention classification achieves 6.5% better accuracy than the Crimson SDK, which suggests that the combination of wavelet packet decomposition along with entropy features are effective in capturing attention state dynamics. Although baseline methods achieve satisfactory results, the data collection strategies of other baseline studies such as Rehman et al. and Huang et al. are based on instructions and there was room for data validation strategies which are implemented in this research. The steps followed in this research ensure correct predictions which pertain to ground truths and this is crucial for real-time implementation.
@@
Initially, device selection was emphasized, as most previous research relied on non-consumer-grade devices. For this study, the FocusCalm headband was chosen for its portability, wearability, and ease of hands-on use, despite the potential for experiments with the Muse headband, which offers greater spatial coverage in real-time applications. Given our objective of developing a state-of-the-art real-time binary classification system, the FocusCalm device has proven sufficient for our purposes. The data collection protocol was designed to reflect the environmental setting of classrooms, with video stimuli consisting of lecture content prepared for students. The data validation process was established under the principle that only ground-truth labels are required. Instances can occur where a participant experiences cognitive load yet fails to answer a question. However, this did not affect labeling, as the protocol ensures that all selected data correspond to verified ground truths.
@@
The methods were selected through extensive trials. For instance, both causal forward filtering and zero-phase second-order section (SOS) filtering were tested. While zero-phase SOS filtering was more effective at removing phase distortion, its non-causal nature performed poorly at segment edges. To address this, those edge portions were removed, ensuring both the absence of distortion and the independence from future data. Furthermore, to prevent data loss, the segmentation process used an overlapping ratio. For classification, SVM was initially selected due to its proven effectiveness in EEG signal analysis. Conceptually, SVM is viewed as a mechanism that iteratively identifies the most discriminative features in a high-dimensional space, analogous to how traditional EEG analyses iteratively isolate regions or frequency bands of interest to distinguish cognitive states. An initial assessment using various machine learning models was conducted, and the SVM achieved the highest performance, which is consistent with observations reported in the literature.
@@
After validating the data using LOSO, the optimal $(C_{2},\gamma)$ values were determined in two ways. The first method identified the most common $(C_{2},\gamma)$ pair that achieved the best metrics across all subjects. The second method selected the $(C_{2},\gamma)$ pair that yielded the highest metrics, regardless of its frequency among subjects. The mode pair, $(C_{2},\gamma)=(0.01,0.5)$, produced an average accuracy of 88.77%, whereas $(C_{2},\gamma)=(0.01,0.1)$ achieved a slightly higher average accuracy of 89.63%. Despite the higher accuracy, the former pair was adopted for real-time testing due to its more frequent favorable performance, which makes the model generalized and objective. The LOSO validation results given in Fig. 4 show that while the model performed well on most subjects, it performed poorly on two. This indicates the model’s robustness could be improved with additional data.
@@
Figure 5: Subfigures (a) and (b) represent the graphical user interface and the real-time working pipeline. Subfigure (c) denotes the changes between attentive and non-attentive states during the real-time system test on five participants.
VI Real-time Setup
VI-A Data-handling Procedure
The real-time system was implemented in Python using a circular buffer for continuous EEG data processing. Fixed-length segments of 1750 samples were extracted with a step size of 525 samples (70% overlap), resulting in a new classification output approximately every 2.1 seconds. Each segment was processed using noise removal, feature extraction, and classification by the trained model.